Quantitative Problem: Barton Industries estimates its cost of common equity by using three approaches: the CAPM, the bond-yield-plus-risk-premium approach, and the DCF model. Barton expects next year's annual dividend, D1, to be $2.10 and it expects dividends to grow at a constant rate g = 5.0%. The firm's current common stock price, P0, is $20.00. The current risk-free rate, rRF, = 4.2%; the market risk premium, RPM, = 5.9%, and the firm's stock has a current beta, b, = 1.30. Assume that the firm's cost of debt, rd, is 16.23%. The firm uses a 2.9% risk premium when arriving at a ballpark estimate of its cost of equity using the bond-yield-plus-risk-premium approach. What is the firm's cost of equity using each of these three approaches? Round your answers to two decimal places.

CAPM cost of equity:   %
Bond yield plus risk premium:   %
DCF cost of equity:   %

What is your best estimate of the firm's cost of equity?
Answer #1

a). According to the CAPM,

Cost of Equity = rRF + [b * RPM] = 4.2% + [1.30 * 5.9%] = 4.2% + 7.67% = 11.87%

b). Cost of Equity = rd + Risk Premium = 16.23% + 2.9% = 19.13%

c). Cost of Equity = [D1 / P0] + g

= [$2.10 / $20] + 0.05

= 0.105 + 0.05 = 0.155, or 15.50%

d). Best estimate will be the average of the three methods

Cost of Equity = [11.87% + 19.13% + 15.50%] / 3 = 46.50% / 3 = 15.50%